OpenStudy (anonymous):

The loudness of the sound in a concert hall and in a football field are 80 dB and 85 dB respectively. Find the ratio of the sound intensities in the football field to the concert hall.

OpenStudy (anonymous):

85db/80db = 16/17 so every 16db in concert hall will have 17db in football field

OpenStudy (kropot72):

The sound intensity on the football field is greater than the sound intensity in the concert hall by (85 - 80)dB.

OpenStudy (ash2326):

Sound Intensity is proportional to the power in watts Relation between decibels and power \[ db= 10 \log_{10} P\] P= power in watts here let power P1 be in concert hall, P2 be the power in the football field Find P2/P1

OpenStudy (ash2326):

@kryton1212 Do you get this?

OpenStudy (anonymous):

yes =] thank you

OpenStudy (ash2326):

@kryton1212 Can you do the next part?

OpenStudy (anonymous):

wait a moment

OpenStudy (ash2326):

Sure

OpenStudy (anonymous):

sorry I cannot get it...

OpenStudy (kropot72):

The sound intensity on the football field is 5dB greater than the sound intensity in the concert hall. The measurements of the two intensities are relative to the same reference level and are logarithmic. Therefore subtraction is required in this case.

OpenStudy (anonymous):

I think the question is asking the ratio not the difference.

OpenStudy (kropot72):

The result of 5dB is a ratio.

OpenStudy (anonymous):

but the answer is 3.2:1

OpenStudy (ash2326):

Yeah, 5 db is ratio. You could convert that to power, that will also be the ratio. \[5=10\log_{10} P\] P is the ratio

OpenStudy (anonymous):

(85-80)dB=10 log P which P is the ratio of the sound intensities

OpenStudy (ash2326):

yes

OpenStudy (kropot72):

A power ratio of 5dB is equivalent to a power ratio of 3.2:1

OpenStudy (anonymous):

why?

OpenStudy (ash2326):

Subtraction of dbs, is equivalent to division of powers. That's why it's a ratio

OpenStudy (anonymous):

5=10logP and then?

OpenStudy (ash2326):

\\[5=10\log_{10} P\] \[\frac{5}{10}=\log_{10} P\] Take anti log both sides \[10^{\frac{5}{10}}=P\] \[10^{\frac 1 2}=P\]

OpenStudy (kropot72):

\[\log_{} P=\frac{5}{10}\] \[10^{0.5}=3.16\]

OpenStudy (anonymous):

thank you. I know now